Ask your own question, for FREE!
Mathematics 9 Online
OpenStudy (anonymous):

There are 52 cards in a deck. There are 4 eights and 4 nines. What is the probability that if you cut the deck, you'll get an eight or a nine? Express your answer as a decimal number rounded to the hundredths place

OpenStudy (anonymous):

4/52 + 4/52= 8/52= 2/13. change to decimal number. is it going right?

OpenStudy (anonymous):

Idk how to change that to adecimal number.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2%2F13

OpenStudy (anonymous):

0.15?

OpenStudy (anonymous):

yes..

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!